OpenStudy (anonymous):

What is the solution to the equation 0=10k^2-8k+8 in terms of k?

OpenStudy (anonymous):

divide both side by 2 - just to simplify. use quadratic formula: k=2+/- 4i

OpenStudy (anonymous):

Use the quadratic equation to solve. \[k = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Where.... \[0 = ak^2 - bk + c\] a, b, and c are the constants in the equation. Substitute in and watch the magic happen. If you still need help, ask me to clarify.

OpenStudy (anonymous):

8 +/- 16i k = ---------- 20 4(2 +/- 4i) 2 +/- 4i k = ---------- = --------- 4*5 5

OpenStudy (llort):

Possible intermediate steps: 0 = 10 k^2-8 k+8 Subtract (10 k^2-8 k+8) from both sides: -10 k^2+8 k-8 = 0 Solve the quadratic equation by completing the square: Divide both sides by -10: k^2-(4 k)/5+4/5 = 0 Subtract 4/5 from both sides: k^2-(4 k)/5 = -4/5 Add 4/25 to both sides: k^2-(4 k)/5+4/25 = -16/25 Factor the left hand side: (k-2/5)^2 = -16/25 Take the square root of both sides: abs(k-2/5) = (4 i)/5 Eliminate the absolute value: k-2/5 = -(4 i)/5 or k-2/5 = (4 i)/5 Add 2/5 to both sides: k = 2/5-(4 i)/5 or k-2/5 = (4 i)/5 Add 2/5 to both sides: k = 2/5-(4 i)/5 or k = 2/5+(4 i)/5

OpenStudy (anonymous):

Oh ok i get it now thank you for explaining.